Two angles are vertical. One angle measures (x+16) degrees and the other
angle measures (4x-5) degrees. Determine the measures of both angles.

Answer by ikleyn(52802)

.
Vertical angles are congruent and have equal measures. Therefore,


x + 16 = 4x -5.


Simplify and solve for x :


16 + 5 = 4x - x

21 = 3x  =================>  x = 21/3 = 7.


Thus the angles measure is  x+16 = 7+16 = 23 degrees, for both angles (for each of the two vertical angles).   ANSWER

Solved.